# Improving Long-Range Navigation with Spatially-Enhanced Recurrent Memory via End-to-End Reinforcement Learning

### v1

@@ Abstract
Recent advancements in robot navigation, particularly with end-to-end learning approaches such as reinforcement learning (RL), have demonstrated remarkable efficiency and effectiveness. However, successful navigation still depends on two key capabilities: mapping and planning, whether implemented explicitly or implicitly. Classical approaches rely on explicit mapping pipelines to transform and register ego-centric observations into a coherent map frame for use by the planner module. In contrast, end-to-end learning can achieve this implicitly, often through recurrent neural networks (RNNs) that fuse current and historical observations into a latent space for planning. While existing architectures, such as LSTM and GRU, can capture temporal dependencies, our findings reveal a critical limitation: their inability to effectively perform spatial memorization. This capability is essential for transforming and integrating sequential observations from varying perspectives to build spatial representations that support downstream planning tasks. To address this, we propose Spatially-Enhanced Recurrent Units (SRUs), a simple yet effective modification to existing RNNs, designed to enhance spatial memorization capabilities. To improve navigation performance, we introduce an attention-based network architecture integrated with SRUs, enabling long-range navigation using a single forward-facing stereo camera. Additionally, we employ regularization techniques to facilitate robust end-to-end recurrent training via RL. Experimental results demonstrate that our approach improves long-range navigation performance by an overall of 23.5% compared to existing RNNs. Furthermore, when equipped with SRU memory, our method outperforms the RL baseline approach, which relies on explicit mapping and memory modules, achieving an overall 29.6% improvement in diverse environments that require long-horizon mapping and memorization capabilities. Finally, we address the sim-to-real gap by leveraging large-scale pretraining on synthetic depth data, enabling zero-shot transfer for deployment across diverse and complex real-world environments.
@@
Spatial Memory, End-to-End Navigation, Recurrent Neural Networks, Reinforcement Learning
@@ 1 Introduction
To address this limitation, we introduce Spatially-Enhanced Recurrent Units (SRUs), a simple yet effective modification to standard LSTM and GRU units that enhances their spatial registration capabilities when processing sequences of ego-centric observations. Unlike classical mapping pipelines that rely on explicit homogeneous transformations, our approach enables the recurrent units to implicitly learn the transformations from varying observation perspectives effectively. To further enhance the performance of long-range navigation tasks, we propose an attention-based network architecture integrated with SRUs, allowing the model to learn long-range navigation policies using only ego-centric observations via end-to-end RL. Our experiments demonstrate improvements in spatial awareness compared to the baselines. With the SRU memory, the implicit recurrent approach via RL with sparse rewards promotes robust exploration in complex 3D and maze-like environments, outperforming the baseline that rely on explicit mapping and memory modules.
@@
  • End-to-End Reinforcement Learning with SRUs and Attention-based Policy: We integrate the SRU unit into a proposed attention-based network architecture, enabling improved end-to-end reinforcement learning for long-range navigation tasks using only ego-centric observations.
  • Large-Scale Pretraining for Zero-Shot Sim-to-Real Transfer in Long-Range Navigation: By leveraging large-scale synthetic pretraining and a parallelizable depth-noise model, our system bridges the sim-to-real gap, enabling zero-shot deployment on a legged-wheel platform in diverse real-world environments, using a single forward-facing stereo camera for long-range navigation.
@@ 2 Related Works
To address these limitations, recent research has increasingly turned to learning-based approaches, especially for more complex robotic agents (e.g., quadrupeds or wheel-legged systems). For instance, recent works in imitation learning leverage large-scale video data or demonstrations (Pfeiffer et al., 2017; Bojarski et al., 2016; Loquercio et al., 2021; Shah et al., 2022, 2023) to directly map raw egocentric sensory inputs to navigation actions. Given the challenges of capturing dynamic, closed-loop interactions from purely offline data, researchers have also explored model-free reinforcement learning (RL) methods (Shi et al., 2019; Wijmans et al., 2019; Hoeller et al., 2021; Truong et al., 2021; Ruiz-Serra et al., 2022; Fu et al., 2022; Choi et al., 2019; Bhattacharya et al., 2024; Lee et al., 2024) that train navigation policies end-to-end by simulating the entire robot dynamics. By replacing the traditional perception, mapping, and planning pipeline with a tailored network—such as architectures based on recurrent networks (Wijmans et al., 2019; Choi et al., 2019; Hoeller et al., 2021) or Transformers with attention mechanisms (Ruiz-Serra et al., 2022; Bhattacharya et al., 2024; Zeng et al., 2024)—these approaches have achieved improvements in navigation tasks as well as in robotic locomotion (Miki et al., 2022a; Yang et al., 2022b; Kareer et al., 2023).
@@
A key challenge with end-to-end approaches is learning a robust state representation from the partial observations provided by egocentric sensors. Recent studies have attempted to mitigate this challenge by incorporating explicit mapping and memory mechanisms (Savinov et al., 2018; Fu et al., 2022; Lee et al., 2024) or by employing specialized network architectures like RNNs (Hoeller et al., 2021; Wijmans et al., 2019; Choi et al., 2019). However, RNNs—originally designed to capture temporal sequences in language tasks (Cho et al., 2014)—are not inherently well-suited for spatial mapping, particularly when processing sequential egocentric observations from continuously changing perspectives. For instance, while prior studies in indoor navigation have shown that spatial cues can be decoded from RNN memories, this effect is limited to employing binary contact sensing and does not extend to high-dimensional visual inputs (Wijmans et al., 2023). Moreover, recent findings suggest that variations in recurrent network architectures have minimal impact on the final task-level rewards achieved through reinforcement learning (Duarte et al., 2023). This indicates that, despite architectural differences, some fundamental limitations may persist across these RNN units.
@@
In this paper, we explore a key limitation of existing RNN-based architectures in addressing partial observability—their spatial memorization capabilities—highlighting their shortcomings in learning spatial transformations and integrating observations from different perspectives. We then introduce Spatially-Enhanced Recurrent Units (SRUs) and demonstrate their effectiveness in improving long-range navigation tasks with a specifically designed attention-based network structure via end-to-end reinforcement learning.
@@ 3 Problem Statement
  • $\gamma\in[0,1]$: the discount factor that balances immediate and future rewards.
@@
Due to the robot’s ego-motion, the current observation $o_{t}$ can be captured from a different perspective or observation frame compared to historical observations in $\mathcal{H}_{t-1}$. Therefore, to fuse observations from different viewpoints, the function $f$ typically involves a spatial transformation that aligns the observations into a coherent reference frame to estimate the current state $s_{t}$ for the policy $\pi$. In classical mapping pipelines, spatial transformations are typically achieved through homogeneous transformations that combine rotations and translations. However, in end-to-end learning approaches, the function $f$ is learned implicitly.
@@ 4.1 Overview
To tackle the long-range navigation task, we first examine and demonstrate the limitations of existing recurrent architectures (e.g., LSTM, GRU, S4, and Mamba-SSM) in a spatial-temporal memory task. We then introduce the Spatially-Enhanced Recurrent Units (SRUs). Next, we integrate SRUs into an attention-based network architecture to learn long-range navigation via end-to-end reinforcement learning. Furthermore, we discuss the importance of incorporating regularization techniques to prevent early overfitting, which we find to be crucial to enhance SRUs’ spatial memorization. Finally, we address the sim-to-real gap by pretraining the depth image encoder on large-scale synthetic depth data and incorporating a parallelizable depth-noise model, enabling zero-shot transfer to real-world environments.
@@ 4.3 Spatial Mapping Limitations in RNNs
Achieving long-range navigation from ego-centric observations requires the robot to perform effective spatial mapping. In three-dimensional (3D) space, spatial mapping is commonly achieved using homogeneous transformations, which combine rotations and translations. A general representation of such a transformation is expressed as:
@@
In this section, we assess the spatial and temporal mapping performance of existing recurrent structures—namely LSTM, GRU, and the recent S4 and Mamba-SSM—on two fronts: (i) temporal memorization and (ii) spatial transformation and memorization. Consider an abstract scenario relevant to the navigation task, in which a robot is initialized at a pose in $SE(3)$ and moves randomly within the three-dimensional environment $E\subset\mathbb{R}^{3}$. At each time step $t$, the robot receives an observation $o_{t}$ containing the coordinates of observed landmarks $l^{i}_{t}\in\mathbb{R}^{3}$, defined relative to the robot’s current frame (indicated by the subscript $t$). Each landmark is also associated with a binary categorical label $c^{i}$, which is temporally relevant and independent of the observation frame. Additionally, the robot is provided with its ego-motion transformation matrix $M^{t-1}_{t}$, representing the transformation from the previous pose at time step $t-1$ to the pose at time step $t$. This transformation enables the network to align and integrate observations into a unified reference frame, akin to classical homogeneous transformations. Over a sequence of $T$ time steps, the RNN processes these observations. At the final step $T$, the network is evaluated simultaneously on its ability to:
@@
Figure 2: Spatial Mapping Comparison: (a) and (b) depict the spatial mapping performance of LSTM and SRU-LSTM units on synthetic data, respectively, as the robot follows a spiral path, observing landmarks from different perspectives. At the end of the path, the robot is tasked with memorizing and transforming the observed landmark coordinates into the final robot frame. Numbers indicate observation time steps. (c) illustrates spatial memory errors across observation step indices, from the final (15) to the initial step (1), using various randomly generated trajectories and observations.
@@ 4.5 Attention-Based Recurrent Network Architecture for Navigation
To leverage the SRUs in the navigation context, we propose an attention-based recurrent network architecture for long-range navigation tasks using raw front-facing stereo depth input, as illustrated in Figure 3. The network consists of a pretrained depth image encoder, two spatial attention layers incorporating both self-attention and cross-attention mechanisms to enhance and compress the encoded visual features, and a recurrent unit (SRU) that learns a spatial-temporal representation of the state by fusing the current observation with historical observations. Finally, a multilayer perceptron (MLP) head computes the actions from the recurrent hidden state, outputting velocity commands for the robot’s locomotion controller.
@@ 4.5.1 Depth Encoder Pretraining and Simulated Perception Noise
For the depth encoder, we adopt a convolutional neural network (CNN) backbone based on RegNet (Radosavovic et al., 2020), enhanced with a Feature Pyramid Network (FPN) (Lin et al., 2017) to capture spatial features across multiple scales. The encoder is pretrained for self-reconstruction on large-scale synthetic depth image data from TartanAir (Wang et al., 2020) using a variational autoencoder (VAE) framework. This pretraining enables the encoder to learn and extract robust and generalizable features from depth images, facilitating effective downstream navigation learning and deployment. However, depth images captured in simulation often differ from those obtained in real-world environments due to various sensor artifacts and noise. To address this sim-to-real gap, we integrate a parallelized depth-noise model, adapted from (Handa et al., 2014b; Barron and Malik, 2013b; Bohg et al., 2014b), which introduces configurable noise to the depth images, such as:
@@ 4.5.2 Attention Layers for Feature Compression
^$\hat{\mathcal{F}}_{t}\in\mathbb{R}^{C\times 1},$
which preserves the most relevant spatial features while reducing the dimensionality. This compressed feature ^$\hat{\mathcal{F}}_{t}$ is then concatenated with the proprioceptive state $o_{t}^{\text{prop}}$ and the relative goal position $p_{t}$ before being passed to the recurrent memory unit. There, it is fused with historical observations to form a spatial-temporal representation of the robot’s current state, integrating both exteroceptive and proprioceptive information.
@@ 4.5.3 Spatially-Enhanced Recurrent Unit (SRU)
The Spatially-Enhanced Recurrent Unit (SRU), as described in Sec. 4.4, processes the compressed feature map ^$\hat{\mathcal{F}}_{t}$, along with the robot’s current proprioceptive state $o_{t}^{\text{prop}}$ and the previous hidden state $h_{t-1}$, to generate a spatial-temporal representation of the surrounding environment from the robot’s egocentric observations. The observation of the current proprioceptive state $o_{t}^{\text{prop}}$ provides essential ego-motion information, equivalent to the transformation matrix $M^{t-1}_{t}$ in Sec. 4.3. The SRU learns to implicitly perform spatial transformations, aligning the current observation feature map ^$\hat{\mathcal{F}}_{t}$ with the previous hidden state $h_{t-1}$. The resulting hidden state $h_{t}$, which encapsulates the integrated environmental information to estimate the robot’s current state $s_{t}$, is subsequently passed through a multi-layer perceptron (MLP) head to compute the robot’s action $a_{t}$.
@@ 4.6 Learning Navigation with Sparse Rewards and Regularizations
The final attention-based network with the SRU is trained end-to-end using RL to achieve long-range navigation, with the objective of maximizing the cumulative reward over the episode. The reward function for the navigation task is designed as a combination of task-level rewards $r^{\text{task}}$, regularization $r^{\text{reg}}$, and penalty $r^{\text{pen}}$ terms, as follows:
@@
The regularization term $r^{\text{reg}}_{t}$ encourages smooth and energy-efficient behaviors by penalizing rapid action changes and excessive joint accelerations. This is implemented using L1 regularization on the difference between the current action $a_{t}$ and a momentum-filtered version of previous actions $a_{t}^{m}$, as defined below:

### v2

# Spatially-Enhanced Recurrent Memory for Long-Range Mapless Navigation via End-to-End Reinforcement Learning

@@ Abstract
Recent advancements in robot navigation, particularly with end-to-end learning approaches such as reinforcement learning (RL), have demonstrated remarkable efficiency and effectiveness. However, successful navigation still depends on two key capabilities: mapping and planning, whether implemented explicitly or implicitly. Classical approaches rely on explicit mapping pipelines to transform and register egocentric observations into a coherent map for the planning module. In contrast, end-to-end learning often achieves this implicitly—through recurrent neural networks (RNNs) that fuse current and historical observations into a latent space for planning. While existing architectures, such as LSTM and GRU, can capture temporal dependencies, our findings reveal a critical limitation: their inability to effectively perform spatial memorization. This capability is essential for transforming and integrating sequential observations from varying perspectives to build spatial representations that support planning tasks. To address this, we propose Spatially-Enhanced Recurrent Units (SRUs)—a simple yet effective modification to existing RNNs—that enhance spatial memorization. To improve navigation performance, we introduce an attention-based network architecture integrated with SRUs, enabling long-range mapless navigation using a single forward-facing stereo camera. Additionally, we employ regularization techniques to facilitate robust end-to-end recurrent training via RL. Experimental results demonstrate that our approach improves long-range navigation performance by 23.5% overall compared to existing RNNs. Furthermore, when equipped with SRU memory, our method outperforms both RL baseline approaches—one relying on explicit mapping and the other on stacked historical observations—achieving overall improvements of 29.6% and 105.0%, respectively, in diverse environments that require long-horizon mapping and memorization capabilities. Finally, we address the sim-to-real gap by leveraging large-scale pretraining on synthetic depth data, enabling zero-shot transfer for deployment across diverse and complex real-world environments.
@@
Spatial Memory, End-to-End Mapless Navigation, Recurrent Neural Networks, Reinforcement Learning
@@ 1 Introduction
To address this limitation, we introduce Spatially-Enhanced Recurrent Units (SRUs), a simple yet effective modification to standard LSTM and GRU units that enhances their spatial registration capabilities when processing sequences of ego-centric observations. Unlike classical mapping pipelines that rely on explicit homogeneous transformations, our approach enables the recurrent units to implicitly learn the transformations from varying observation perspectives effectively. To further enhance the performance of long-range navigation tasks, we propose an attention-based network architecture integrated with SRUs, allowing the model to learn long-range mapless navigation policies using only ego-centric observations via end-to-end RL. Our experiments demonstrate improvements in spatial awareness compared to the baselines. With the SRU memory, the implicit recurrent approach via RL with sparse rewards promotes robust exploration in complex 3D and maze-like environments, outperforming the baseline that rely on explicit mapping and memory modules.
@@
  • End-to-End Reinforcement Learning with SRUs and Attention-based Policy: We integrate the SRU unit into a proposed attention-based network architecture, enabling improved end-to-end reinforcement learning for long-range mapless navigation tasks using only ego-centric observations.
  • Large-Scale Pretraining for Zero-Shot Sim-to-Real Transfer in Long-Range Mapless Navigation: By leveraging large-scale synthetic pretraining and a parallelizable depth-noise model, our system bridges the sim-to-real gap, enabling zero-shot deployment on a legged-wheel platform in diverse real-world environments, using a single forward-facing stereo camera for long-range mapless navigation.
@@ 2 Related Works
To address these limitations, recent research has increasingly turned to learning-based approaches, especially for more complex robotic agents (e.g., quadrupeds or wheel-legged systems). For instance, recent works in imitation learning leverage large-scale video data or demonstrations (Pfeiffer et al., 2017; Bojarski et al., 2016; Loquercio et al., 2021; Shah et al., 2022, 2023) to directly map raw egocentric sensory inputs to navigation actions. Given the challenges of capturing dynamic, closed-loop interactions from purely offline data, researchers have also explored model-free reinforcement learning (RL) methods (Shi et al., 2019; Wijmans et al., 2019; Choi et al., 2019; Hoeller et al., 2021; Truong et al., 2021; Wu et al., 2021; Ruiz-Serra et al., 2022; Fu et al., 2022; Huang et al., 2023; Bhattacharya et al., 2024; Lee et al., 2024) that train navigation policies end-to-end by simulating the entire robot dynamics. By replacing the traditional perception, mapping, and planning pipeline with a tailored network—such as architectures based on recurrent networks (Wijmans et al., 2019; Choi et al., 2019; Hoeller et al., 2021; Wu et al., 2021) or Transformers with attention mechanisms (Ruiz-Serra et al., 2022; Huang et al., 2023; Bhattacharya et al., 2024; Zeng et al., 2024)—these approaches have achieved improvements in navigation tasks as well as in robotic locomotion (Miki et al., 2022a; Yang et al., 2022b; Kareer et al., 2023).
@@
A key challenge with end-to-end approaches is learning a robust state representation from the partial observations provided by egocentric sensors. Recent studies have attempted to mitigate this challenge by incorporating explicit mapping and memory mechanisms (Savinov et al., 2018; Cimurs et al., 2021; Fu et al., 2022; Lee et al., 2024) or by employing specialized network architectures like RNNs (Hoeller et al., 2021; Wijmans et al., 2019; Choi et al., 2019). However, RNNs—originally designed to capture temporal sequences in language tasks (Cho et al., 2014)—are not inherently well-suited for spatial mapping, particularly when processing sequential egocentric observations from continuously changing perspectives. For instance, while prior studies in indoor navigation have shown that spatial cues can be decoded from RNN memories, this effect is limited to employing binary contact sensing and does not extend to high-dimensional visual inputs (Wijmans et al., 2023). Moreover, recent findings suggest that variations in recurrent network architectures have minimal impact on the final task-level rewards achieved through reinforcement learning (Duarte et al., 2023). This indicates that, despite architectural differences, some fundamental limitations may persist across these RNN units.
@@
In this paper, we explore a key limitation of existing RNN-based architectures in addressing partial observability—their spatial memorization capabilities—highlighting their shortcomings in learning spatial transformations and integrating observations from different perspectives. We then introduce Spatially-Enhanced Recurrent Units (SRUs) and demonstrate their effectiveness in improving long-range mapless navigation tasks with a specifically designed attention-based network structure via end-to-end reinforcement learning.
@@ 3 Problem Statement
  • $\gamma\in[0,1]$: the discount factor that balances immediate reward and future payoffs.
@@
Due to the robot’s ego-motion, the current observation $o_{t}$ can be captured from a different perspective or observation frame compared to historical observations in $\mathcal{H}_{t-1}$. Therefore, to fuse observations from different viewpoints, the function $f$ typically involves a spatial transformation that aligns the observations into a coherent reference frame to estimate the current state $s_{t}$ for the policy $\pi$. In classical mapping pipelines, spatial transformations are typically achieved through homogeneous transformations that combine rotations and translations. However, in end-to-end learning approaches, the function $f$, which can be parameterized by neural network weights, is learned implicitly.
@@ 4.1 Overview
To tackle the long-range navigation task, we first examine and demonstrate the limitations of existing recurrent architectures (e.g., LSTM, GRU, S4, and Mamba-SSM) in a spatial-temporal memory task. We then introduce the Spatially-Enhanced Recurrent Units (SRUs). Next, we integrate SRUs into an attention-based network architecture to learn long-range mapless navigation via end-to-end reinforcement learning. Furthermore, we discuss the importance of incorporating regularization techniques to prevent early overfitting, which we find to be crucial to enhance SRUs’ spatial memorization. Finally, we address the sim-to-real gap by pretraining the depth image encoder on large-scale synthetic depth data and incorporating a parallelizable depth-noise model, enabling zero-shot transfer to real-world environments.
@@ 4.3 Spatial Mapping Limitations in RNNs
Achieving long-range mapless navigation from ego-centric observations requires the robot to perform effective spatial mapping. In three-dimensional (3D) space, spatial mapping is commonly achieved using homogeneous transformations, which combine rotations and translations. A general representation of such a transformation is expressed as:
@@
In this section, we assess the spatial and temporal mapping performance of existing recurrent structures—namely LSTM, GRU, and the recent S4 and Mamba-SSM—on two fronts: (i) temporal memorization and (ii) spatial transformation and memorization. Consider an abstract scenario relevant to the navigation task, in which a robot is initialized at a pose in $SE(3)$ and moves randomly within the three-dimensional environment $E\subset\mathbb{R}^{3}$. At each time step $t$, the robot receives an observation $o_{t}$ containing the coordinates of observed landmarks $l^{i}_{t}\in\mathbb{R}^{3}$, defined relative to the robot’s current frame (indicated by the subscript $t$). Each landmark is also associated with a binary categorical label $c^{i}$, which is temporally relevant and independent of the observation frame. Additionally, the robot is provided with its ego-motion transformation matrix $M^{t-1}_{t}$, representing the transformation from the previous pose at time step t1t-1 to the pose at time step $t$. This transformation enables the network to align and integrate observations into a unified reference frame, akin to classical homogeneous transformations. Over a sequence of $T$ time steps, the RNN processes these observations. At the final step $T$, the network is evaluated simultaneously on its ability to:
@@
Figure 2: Spatial Mapping Comparison: (a) and (b) depict the spatial mapping performance of LSTM and SRU-LSTM units on synthetic data, respectively, as the robot follows a spiral path, observing landmarks from different perspectives. At the end of the path, the robot is tasked with memorizing and transforming the observed landmark coordinates into the final robot frame. Numbers indicate observation time steps. (c) illustrates the mean spatial memory errors (log scale) across observation step indices, ordered from the final (15) to the initial step (1), averaged over various randomly generated trajectories and observations.
@@ 4.5 Attention-Based Recurrent Network Architecture for Navigation
To leverage the SRUs in the navigation context, we propose an attention-based recurrent network architecture for long-range mapless navigation tasks using raw front-facing stereo depth input, as illustrated in Figure 3. The network consists of a pretrained depth image encoder, two spatial attention layers incorporating both self-attention and cross-attention mechanisms to enhance and compress the encoded visual features, and a recurrent unit (SRU) that learns a spatial-temporal representation of the state by fusing the current observation with historical observations. Finally, a multilayer perceptron (MLP) head computes the actions from the recurrent hidden state, outputting velocity commands for the robot’s locomotion controller.
@@ 4.5.1 Depth Encoder Pretraining and Simulated Perception Noise
For the depth encoder, we adopt a convolutional neural network (CNN) backbone based on RegNet (Radosavovic et al., 2020), chosen for its simplicity and efficiency. This is further enhanced with a Feature Pyramid Network (FPN) (Lin et al., 2017) to capture spatial features across multiple scales. The encoder is pretrained for self-reconstruction on large-scale synthetic depth image data from TartanAir (Wang et al., 2020) using a variational autoencoder (VAE) framework. This pretraining enables the encoder to learn and extract robust and generalizable features from depth images, facilitating effective downstream navigation learning and deployment. However, depth images captured in simulation often differ from those obtained in real-world environments due to various sensor artifacts and noise. To address this sim-to-real gap, we integrate a parallelized depth-noise model, adapted from (Handa et al., 2014b; Barron and Malik, 2013b; Bohg et al., 2014b), which introduces configurable noise to the depth images, such as:
@@ 4.5.2 Attention Layers for Feature Compression
$\hat{\mathcal{F}}_{t}\in\mathbb{R}^{C\times 1},$
which preserves the most relevant spatial features while reducing the dimensionality. This compressed feature $\hat{\mathcal{F}}_{t}$ is then concatenated with the proprioceptive state $o_{t}^{\text{prop}}$ and the relative goal position $p_{t}$ before being passed to the recurrent memory unit. There, it is fused with historical observations to form a spatial-temporal representation of the robot’s current state, integrating both exteroceptive and proprioceptive information.
@@ 4.5.3 Spatially-Enhanced Recurrent Unit (SRU)
The Spatially-Enhanced Recurrent Unit (SRU), as described in Sec. 4.4, processes the compressed feature map $\hat{\mathcal{F}}_{t}$, along with the robot’s current proprioceptive state $o_{t}^{\text{prop}}$ and the previous hidden state $h_{t-1}$, to generate a spatial-temporal representation of the surrounding environment from the robot’s egocentric observations. The observation of the current proprioceptive state $o_{t}^{\text{prop}}$ provides essential ego-motion information, equivalent to the transformation matrix $M^{t-1}_{t}$ in Sec. 4.3. The SRU learns to implicitly perform spatial transformations, aligning the current observation feature map $\hat{\mathcal{F}}_{t}$ with the previous hidden state $h_{t-1}$. The resulting hidden state $h_{t}$, which encapsulates the integrated environmental information to estimate the robot’s current state $s_{t}$, is subsequently passed through a multi-layer perceptron (MLP) head to compute the robot’s action $a_{t}$.
@@ 4.6 Learning Navigation with Sparse Rewards and Regularizations
The final attention-based network with the SRU is trained end-to-end using RL to achieve long-range mapless navigation, with the objective of maximizing the cumulative reward over the episode. The reward function for the navigation task is designed as a combination of task-level rewards $r^{\text{task}}$, regularization $r^{\text{reg}}$, and penalty $r^{\text{pen}}$ terms, as follows:
@@
The regularization term $r^{\text{reg}}_{t}$ encourages smooth behaviors by penalizing rapid action changes and excessive joint accelerations. This is implemented using L1 regularization on the difference between the current action $a_{t}$ and a momentum-filtered version of previous actions $a_{t}^{m}$, as defined below: